# Learning to Deblur Polarized Images

### v1

@@ 6.2 Evaluation on real data
To show that our method has a good generalization ability on real images, we capture several blurry images from various scenes using a Lucid Vision Phoenix polarization camera. Qualitative results are shown in Fig. $4$Additional results can be found in the supplementary material.. From the results we can see that our method suffer less from artifacts. Taking the left box of $\mathbf{p}$ in the first group of images as an example, the compared methods Cho et al (2021); Ji et al (2022); Zhang et al (2022a); Zhou et al (2023b) tend to produce inaccurate results (shown as abnormal colors). Taking the right box of $\mathbf{I}$ in the first group of images as another example, the compared methods Cho et al (2021); Ji et al (2022); Zhang et al (2022a); Zhou et al (2023b) fail to restore sharp image content.
Figure 5: Results of polarization-based vision applications (including (a) image dehazing (Zhou et al, 2021) and (b) reflection removal (Lyu et al, 2019, 2022)) without deblurring and deblurred by our method and the compared methods (MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and PLIE (Zhou et al, 2023b)). Please zoom-in for better details.
@@ 6.3 Ablation study
    W/o $\hat{\mathbf{S}}_{1,2}$: removing $\hat{\mathbf{S}}_{1,2}$ from $g_{1}$ (to verify the necessity of using $\hat{\mathbf{S}}_{1,2}$ as priors).
@@
    $\mathbf{G}$ instead of $\hat{\mathbf{S}}_{1,2}$: substituting the Stokes parameters $\hat{\mathbf{S}}_{1,2}$ with the gradient distribution $\mathbf{G}$ (to show the advantage of using $\hat{\mathbf{S}}_{1,2}$ over using $\mathbf{G}$).
@@ 6.4 Applications
To show the benefits of deblurring polarized images, we choose two typical polarization-based vision applications including image dehazing (Zhou et al, 2021) and reflection removal (Lyu et al, 2019, 2022) as the validation methods respectively, and show that deblurring can improve their performance. Comparisons of the haze-removed and reflection-removed unpolarized images are shown in Fig. 5 (a) and (b) respectively [^5] Additional results can be found in the supplementary material., which demonstrate that our method brings superior performance gain than the compared methods (Cho et al, 2021; Ji et al, 2022; Zhang et al, 2022a; Zhou et al, 2023b). Taking the red box in Fig. 5 (a) as an example, we can see that the haze cannot be removed robustly without deblurring, and the results deblurred by the compared methods are still hazy, while the result deblurred by our method has higher clarity. Taking the green box in Fig. 5 (b) as another example, we can see that the result deblurred by our method has sharper textures with less reflection contamination than the input and the results deblurred by the compared methods.
@@ 7 Conclusion
We present a learning-based solution to deblur multiple polarized images for acquiring the DoP and AoP accurately. Specifically, we propose a polarized image deblurring pipeline to solve the problem in a polarization-aware manner by decomposing it into two less-ill-posed sub-problems, and build a two-stage neural network tailored to the pipeline to handle the two sub-problems respectively. Experimental results not only show that our method achieves state-of-the-art performance on both synthetic and real-world images, but also demonstrate that using our method to deblur polarized images can improve the performance of polarization-based vision applications such as reflection removal and image dehazing.
Limitations. Since our method is specially designed for deblurring multiple polarized images captured by a polarization camera in a snapshot, it cannot be used to deblur videos since it does not explicitly consider the temporal information encoded in consecutive frames.

### v2

@@ 6.2 Evaluation on real data
To demonstrate the strong generalization capability of our method on complex and diverse real-world data, we capture a real-world dataset using a Lucid Vision Phoenix polarization camera. The dataset comprises 40 different scenes, which exhibits a range of conditions (e.g., different environments, varying blur patterns, etc.). Note that the dataset does not provide the ground truth for reference.
To evaluate the results quantitatively, we choose to adopt a non-reference metric, CLIP-IQA (Wang et al, 2023) (contrastive language-image pretraining for image quality assessment). The results, presented in Table 2, indicate that our model outperforms the compared methods, showcasing superior performance across the dataset. Visual quality comparisons are shown in Fig. $4$Additional results can be found in the supplementary material.. From the results we can see that our method suffer less from artifacts. Taking the first box of $\mathbf{p}$ as an example, the compared methods tend to produce inaccurate results (shown as abnormal colors). Taking the first box of $\mathbf{I}$ as another example, we can see that IPLNet (Hu et al, 2020) brings about ringing artifacts with higher noise level, while other compared methods fail to restore sharp image content.
Table 4: Computational complexity analysis on synthetic data among our method, IPLNet (Hu et al, 2020), ColorPolarNet (Xu et al, 2022), PLIE (Zhou et al, 2023b), MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), STDAN (Zhang et al, 2022a), and MISCFilter (Liu et al, 2024). Here, the size of the input images is $512\times 512$ pixels (3 color channels), and the inference time is the time taken to reconstruct all 300 scenes.
Number of parameters FLOPs Memory Inference time
Ours 3.14 M 360.20 G 2.39 G 4.36 s
IPLNet (Hu et al, 2020) 1.80 M 944.97 G 1.89 G 3.57 s
ColorPolarNet (Xu et al, 2022) 1.03 M 943.94 G 1.51 G 1.90 s
PLIE (Zhou et al, 2023b) 1.50 M 453.94 G 4.28 G 4.84 s
MIMO (Cho et al, 2021) 16.11 M 1207.58 G 2.46 G 13.56 s
XYDeblur (Ji et al, 2022) 4.92 M 550.40 G 2.18 G 12.19 s
STDAN (Zhang et al, 2022a) 13.84 M 3766.04 G 5.94 G 129.73 s
MISCFilter (Liu et al, 2024) 15.99 M 993.29 G 2.82 G 60.92 s
@@ 6.3 Ablation study
    W/o $\mathbf{S}^{*}_{1,2}$: removing $\mathbf{S}^{*}_{1,2}$ from $g_{1}$ (to verify the necessity of using $\mathbf{S}^{*}_{1,2}$ as priors).
@@
    $\mathbf{G}$ instead of $\mathbf{S}^{*}_{1,2}$: substituting the Stokes parameters $\mathbf{S}^{*}_{1,2}$ with the gradient distribution $\mathbf{G}$ (to show the advantage of using $\mathbf{S}^{*}_{1,2}$ over using $\mathbf{G}$).
@@ 6.4 Applications
To show the benefits of deblurring polarized images, we choose two different polarization-based vision applications including image dehazing (Zhou et al, 2021) and reflection removal (Lyu et al, 2019, 2022) as the validation methods respectively, and show that deblurring can improve their performance. Comparisons of the haze-removed and reflection-removed unpolarized images are shown in Fig. 5 (a) and (b) respectively [^5] Additional results can be found in the supplementary material., which demonstrate that our method brings superior performance gain than the compared methods. Taking the red box in Fig. 5 (a) as an example, we can see that the haze cannot be removed robustly without deblurring, and the results deblurred by the compared methods are still hazy, while the result deblurred by our method has higher clarity. Taking the green box in Fig. 5 (b) as another example, we can see that the result deblurred by our method has sharp textures with less reflection contamination, while IPLNet (Hu et al, 2020) alters the contrast of the input image excessively and generates fake textures that originally did not exist, and other compared methods fail to produce sharp textures.
6.5 Computational complexity analysis
To evaluate the computational complexity, we compare the number of parameters, FLOPs, memory, and inference time of different methods on our synthetic test dataset using a single NVIDIA 4090 GPU, as shown in Table 4. Here, the size of the input images is $512\times 512$ pixels (3 color channels), and the inference time is the time taken to reconstruct all 300 scenes. Besides, since the methods designed for handling the unpolarized image deblurring tasks (MIMO (Cho et al, 2021), XYDeblur (Ji et al, 2022), and MISCFilter (Liu et al, 2024)) cannot simultaneously process four polarized images, the total time for processing four polarized images is taken as the inference time. From the results we can see that the computational complexity of our method is not very high, and it can realize real-time deblurring of polarized images.
@@ 7 Conclusion
We present a learning-based solution to deblur multiple polarized images for acquiring the DoLP and AoLP accurately. Specifically, we propose a polarized image deblurring pipeline to solve the problem in a polarization-aware manner by decomposing it into two less-ill-posed sub-problems, and build a two-stage neural network tailored to the pipeline to handle the two sub-problems respectively. Experimental results not only show that our method achieves state-of-the-art performance on both synthetic and real-world images, but also demonstrate that using our method to deblur polarized images can improve the performance of polarization-based vision applications such as reflection removal and image dehazing.
Limitations. Since our method is specially designed for deblurring multiple polarized images captured by a polarization camera in a snapshot, it cannot be used to deblur videos since it does not explicitly consider the temporal information encoded in consecutive frames. Besides, since the formation of motion blur could be different from other types of degeneration (such as haze, rain, and noise), our method, which is based on the observations about motion blur, may not generalize to other low-level tasks.
Discussions. In the future, we aim to extend our work to address the polarized video deblurring problem by developing new modules that can better explore and utilize temporal connections. Additionally, we plan to generalize our approach to handle other types of image degradation by integrating comprehensive observations about polarized images, ultimately working towards a unified framework for enhancing polarized images.
Acknowledgement
This work was supported in part by National Natural Science Foundation of China under Grant No. 62136001, 62088102, 62276007, Beijing Municipal Science & Technology Commission, Administrative Commission of Zhongguancun Science Park (Grant No. Z241100003524012), and JST-Mirai Program Grant Number JPMJM123G1.